OpenStudy (anonymous):

x~lognormal(10,9),compute Pr(3.5<=x<=10).

OpenStudy (anonymous):

You have to use the table for lognormal distribution.

OpenStudy (anonymous):

we have to use the normal table only

OpenStudy (anonymous):

Find Probability [ ln(3.5) < Z < ln(10]=Probability [ 1.25276 < Z <2.30259] Where Z =N[3.5,10]

OpenStudy (anonymous):

Or find \[ P\left[\frac{1}{9} (\ln (3.5)-10)<Z<\frac{1}{9} (\ln (10)-10)\right]\\ =P[-0.971915<Z<-0.855268] \] Where Z=N[0,1]

OpenStudy (anonymous):

P[-0.971915<Z<-0.855268]=0.0306549

OpenStudy (anonymous):

thank you elias,i think there instead 9 is 3

OpenStudy (anonymous):

Sorry, The answer is now 0.00337264

OpenStudy (anonymous):

P(-2.91575<Z<-2.5658)=0.00337264